OpenStudy (anonymous):

If f(x) = 2^(8x^3+1), then find the exact value of f'(0.5)

OpenStudy (anonymous):

Umm, do you know how to find the derivative?

OpenStudy (anonymous):

the derivative of \(b^x\) is \(b^x\ln(b)\) then chain rule

OpenStudy (anonymous):

One way to look at it is: \[ 2^{f(x)} = \left(e^{\ln(2)}\right)^{f(x)} = e^{\ln(2)f(x)} \]Chain rule says: \[ \left(e^{\ln(2)f(x)}\right)' = e^{\ln(2)f(x)}(\ln(2)f(x))' =2^{f(x)}\ln(2)f'(x) \]